# Part I: Nonlinear Codes

### v1

@@ II-B The shift operator ${\bf E}$
A very important operator with whom many of the properties of feedback shift registers can be represented is the shift operator ${\bf E}$. This operator is shifting the whole sequence to the left. For a sequence ${\cal S}=s_{0}s_{1}~{}\cdots~{}s_{k-1}$ of length $k$, the shift operator ${\bf E}s_{i}=s_{i+1}$, where $10\leq i<k-$ 1 and if the sequence is cyclic then also ${\bf E}s_{k-1}=s_{0}$. When the operator is applied on the whole cyclic sequence ${\cal S}$ we have that
@@
Any periodic sequence satisfies a linear recursion
@@
Henceforth, we denote by ${\bf 0}$ any sequence (or a matrix) of any size which contains only zeros. The linear complexity $c({\cal S})$ of ${\cal S}$ is defined as the least $m$ for which (1) holds. Clearly ${c({\cal S})\leq k}$, for any sequence with period $k$, since $s_{i+k}=s_{i}$ by definition.
The sequences that are of interest in our exposition have length which is a power of 2. These sequences are generated by the polynomial $({\bf E}+1)^{k}$ and for each sequence ${\cal S}$ generated by this polynomial we have that $({\bf E}+{\bf 1})^{k}{\cal S}={\bf 0}$. The following properties characterize these sequences [10].
@@
A sequence ${\cal S}$ of length $2^{n}$ has linear complexity $c$ if and only if $({\bf E}+{\bf 1})^{c-1}{\cal S}=1^{2^{n}}$.
@@
There are $2^{n-\lfloor\log n\rfloor-1}$ sequences which are generated by the polynomial $({\bf E}+{\bf 1})^{n+1}$ and are not generated by the polynomial $({\bf E}+{\bf 1})^{n}$. All these sequences have period $2^{\lfloor\log n\rfloor+1}$ and linear complexity ${n+1}$. Each $n$-tuple is contained exactly once in one of these sequences.
We define the weight of a word $(x_{1},x_{2},\ldots,x_{n})$ as the number of nonzero entries in the word. Similarly, we define the weight of a cyclic sequence $[x_{1},x_{2},\ldots,x_{n}]$ as the number of nonzero entries in the sequence. Note, that the weight of the word $(011)$ is two, of the sequence $[011]$ is also two, but of the sequence $[011011]$ is four, although its period is three and the weight of one period is two. Some other properties of sequences can be characterized as follows.
@@ II-C The ${\bf D}$-morphism
A key operator in the recursive constructions which follow is the ${\bf D}$-morphism which was defined by Lempel [27]. The operator ${\bf D}$ is a 2-to-1 mapping from $\mathbb{F}_{2}^{n}$ on to $\mathbb{F}_{2}^{n-1}$, where $n\geq 2$. For a binary word $(x_{1},x_{2},\ldots,x_{n})$ we have
@@
By this definition we have that ${\bf D}x={\bf D}y$ if and only if $y=x$ or $y=\bar{x}$. The operator ${\bf D}$ can be apply on cyclic sequences (cycles). For a cyclic sequence $[x_{1},x_{2},\ldots,x_{n}]$ we have
@@
The definition of ${\bf D}$ implies that ${\bf D}={\bf E}+{\bf 1}$ and hence some of the results concerning the shift operator ${\bf E}$ associated with sequences of length $2^{n}$ can be given in terms of the operator ${\bf D}$, e.g.,
@@
The operator ${\bf D}$ and its inverse ${\bf D}^{-1}$ can be applied several consecutive times on a sequence ${\cal S}$ and the derived sequences. It is defined for $k\geq 2$ as follows.
@@
The equality ${\bf D}={\bf E}+{\bf 1}$ and the properties the operator ${\bf D}$, its inverse ${\bf D}^{-1}$, and the linear complexity imply the following consequences.
@@
If ${\cal F}$ is a set of $\Delta$ vertex-disjoint cycles of length $2^{k}$, , with linear complexity smaller than $2^{k}$, in $G_{n}$, then ${\cal D}^{-1}{\cal F}$ contains $2\Delta$ vertex-disjoint cycles of length $2^{k}$ in $G_{n+1}$.
@@
If ${\cal F}$ is a set of $\Delta$ vertex-disjoint cycles of length $2^{k}$, with linear complexity $2^{k}$, in $G_{n}$, then ${\cal D}^{-1}{\cal F}$ contains $\Delta$ vertex-disjoint self-dual cycles of length $2^{k+1}$ in $G_{n+1}$.
@@ II-D Perfect factors
The most important factors in $G_{n}$ which will play a pivotal role in our exposition are the perfect factors.
@@
The theorem was extended for larger alphabet by Paterson [43]. The following case of his result is required for Theorem 7.
@@
For a given set of parameters there could be different perfect factors. We will be interested in perfect factors in which the weights of all the sequences are either even or odd. Moreover, in these perfect factors all the sequences have the same linear complexity. Both types of factors will be important in our exposition. For sequences of even weight we can characterize all the parameters of perfect factors.
Theorem 4
  1. 1.
  2. 2.
Proof:
  1. 1.
    If $k\leq n<2^{k-1}$, then consider a span $k$ de Bruijn sequence ${\cal S}$ with linear complexity $2^{k-1}+k$ whose existence is implied by Theorem 1. By Lemmas 810, and 11 we have that ${\bf D}^{-r}{\cal S}$ for any $r$ such that $0\leq r<2^{k-1}-k$ contains $2^{r}$ sequences of length $2^{k}$ in $G_{k+r}$, i.e., $2^{n-k}$ sequences of length $2^{k}$ in $G_{n}$ which is a PF$(n,k)$ for any given $n$ such that $k\leq n<2^{k-1}$. By Lemma 8 all the generated sequences have linear complexity $2^{k-1}+k+r<2^{k}$ and hence by Lemma 6 all these sequences have even weight.
  2. 2.
    For $n=2^{k}-1$ let ${\cal F}$ be perfect factor PF$(n,k)$ with sequences of length $n+1=2^{k}$. One of these sequences contains the all-zero word of length $n$. Clearly, the sequence which contains the all-zero word of length $n+1$ is $[0^{n}1]$. This sequence contains all words of length $n$ and weight one. Since each cyclic sequence of length $n+1$ and weight two contains exactly two words of weight one (by omitting a one from the sequence), it follows that ${\cal F}$ does not contain sequences of weight two and hence all words of weight two are contained in sequences with weight three. This implies that ${\cal F}$ does not contain sequences of weight four and so on by induction. Thus, there is a unique factor whose sequences have odd weight. This factor is the state diagram of the CSRn.
Theorem 4 implies that whenever a perfect factor PF$(n,k)$ exists, except for $n=2^{k}-1$, i.e., when $k\leq n<2^{k}-1$ by Theorem 2, there exists such a perfect factor for which all its sequences have even weight. The situation is more complicated for odd weight. PF$(n,n)$ contains exactly one sequence of length $2^{n}$ in $G_{n}$, i.e., a span $n$ de Bruijn sequence. Hence, its weight is always even. For all the other parameters a perfect factor in which all the sequences have odd weight can exist. Unfortunately, whether such perfect factors exist is an open problem. By Theorem 4, if $n=2^{k}-1$, then all the sequences in a PF$(n,k)$ have odd weight. For the other parameters we can use the same technique used in the proof of Theorem 4 and this is left for the reader. In other words we can start with a span $k$ de Bruijn sequence whose linear complexity is $2^{k-1}+\delta$, where $k\leq\delta\leq 2^{k-1}-k-1$ and $\delta\neq k+1$, and apply the inverse operator ${\bf D}^{-r}$, $r>0$, until we reach sequences with linear complexity $2^{n}$, where $n\geq k$. The obtained sequences will be of odd weight by Lemma 6. Other than these two techniques we do not know how to construct perfect factors whose sequences have only odd weights.
It is important to note that in all the constructions that were mentioned, if a sequence ${\cal S}$ is a sequence in the perfect factor ${\cal F}$, then $\bar{{\cal S}}$ is also a sequence in ${\cal F}$, unless ${\cal F}$ contains exactly one sequence. By Lemma 7 all these sequences of length $2^{k}$ are self-dual if and only if they are of linear complexity $12^{k-1}+$ 1. But, it should be noted that other perfect factors without these properties can be also constructed.
Example 2
@@
Finally, one type of PF$(n,n-1)$ is of special interest. In this type of a perfect factor there are two sequences ${\cal S}$ and $\bar{{\cal S}}$. The sequence which contains the all-zero $n$-tuple in such a factor will be called a span $n$ half de Bruijn sequence. In such a sequence either $x$ or $\bar{x}$ is contained in ${\cal S}$, for any given $n$-tuple $x$.
II-E Graphical presentation of two-dimensional arrays
One-dimensional sequences are viewed as cycles in the de Bruijn graph. We aim to view the two-dimensional arrays that will be constructed in our work also as cycles in graphs which will be defined now.
A first interesting and important graph was presented by Fan, Fan, Ma, and Siu [17]. The graph $G_{n,m}(V,E_{1},E_{2})$ has a set of vertices $V$ which consists of all the $2^{nm}$ binary $n\times m$ matrices. There are two sets of edges in the graph, $E_{1}$ and $E_{2}$. The edges in $E_{1}$ are vertical edges and the edges in $E_{2}$ are horizontal edges. Two vertices ($n\times m$ matrices) $v$ and $u$ of $V$ are connected by an edge of $E_{1}$ from $u$ to $v$ if the $(n-1)\times m$ sub-matrix induced by the last $n-1$ rows of $u$ equals to the $(n-1)\times m$ sub-matrix induced by the first $n-1$ rows of $v$. Two vertices ($n\times m$ matrices) $v$ and $u$ of $V$ are connected by an edge of $E_{2}$ from $u$ to $v$ if the $n\times(m-1)$ sub-matrix induced by the last $m-1$ columns of $u$ equals to the $n\times(m-1)$ sub-matrix induced by the first $m-1$ columns of $v$.
A set of $2^{nm-k-t}$ doubly-periodic matrices of size $2^{k}\times 2^{t}$ which is an $(2^{k},2^{t};n,m)$-DBAC $\mathbb{C}$ is equivalent to a set of $2^{nm-k-t}$ disjoint components in the graph $G_{n,m}(V,E_{1},E_{2})$, where each component, associated with a $2^{k}\times 2^{t}$ matrix ${\cal C}\in\mathbb{C}$, contains $2^{k+t}$ vertices which are the $n\times m$ matrices induced by the $2^{k+t}$ windows of size $n\times m$ in ${\cal C}$. These vertices and their connecting edges in one component contain $2^{t}$ vertex-disjoint cycles of length $2^{k}$ with edges from $E_{1}$ and $2^{k}$ vertex disjoint cycles of length $2^{t}$ with edges from $E_{2}$. In other words, they form a factor with edges from $E_{1}$ and a factor with edges from $E_{2}$ in the subgraph induced by the vertices of the component induced by ${\cal C}$. Moreover, it is required that a cycle of length $2^{k}$ with edges from $E_{1}$ is associated with the $2^{k}$ $n\times m$ matrices obtained from a $2^{k}\times m$ sub-matrix of ${\cal C}$. A cycle of length $2^{t}$ with edges from $E_{2}$ is associated with the $2^{t}$ $n\times m$ matrices obtained from an $n\times 2^{t}$ sub-matrix of ${\cal C}$.
@@ II-E Graphical presentation of two-dimensional arrays
The second graph for the representation of the two-dimensional codes was defined before in [8]. This graph is appropriate to codes in which all the columns of the matrices that form the codewords are sequences of some given perfect factor. For this purpose, let ${\cal F}$ be a PF$(n,k)$ and define a graph $G_{{\cal F},m}$ whose vertices are $2^{k}\times m$ matrices. These matrices are periodic only vertically. Each such a matrix $X$ can be represented as
@@
where $X_{i_{\ell}}$, $10\leq\ell\leq m-$ 1, is a cycle in ${\cal F}$ represented as a column vector and ${\bf E}^{j_{\ell}}X_{i_{\ell}}$ is a cyclic shift up of $X_{i_{\ell}}$ by $j_{\ell}$ positions compared to its zero state, i.e., the column vector of ${\bf E}^{j_{\ell}}X_{i_{\ell}}$. The first cycle $X_{i_{0}}$ is taken in its zero shift. An edge in the graph is a $2^{k}\times(m+1)$ matrix which is periodic only vertically. Such a matrix
@@
The graph $G_{{\cal F},m}$ has $2^{nm-k}$ vertices and $2^{n(m+1)-k}$ edges. An Hamiltonian cycle in $G_{{\cal F},m}$ can be represented by a vertically periodic $2^{k}\times(2^{nm-k}+m-1)$ matrix in which each $n\times m$ matrix is contained exactly once. If the $2^{k}\times(m-1)$ matrix obtained from the first $m-1$ columns of this matrix equals (without a cyclic shift) to the $2^{k}\times(m-1)$ matrix obtained from the last $m-1$ columns, then the Hamiltonian cycle can be represented by a $2^{k}\times 2^{nm-k}$ doubly-periodic matrix. This condition is satisfied when $m>2$ or $n\neq k$ and the associated matrix forms a $(2^{k},2^{mn-k};n,m)$-DBAC with exactly one codeword, i.e., a $(2^{k},2^{mn-k};n,m)$-DBA. When $m=2$ and $n=k$ the last column will be shifted in $2^{n-1}$ positions compared to the first one and we cannot remove it to form a doubly-periodic array. This is something that has to be taken into account in our constructions too.
Similarly, an Eulerian cycle in $G_{{\cal F},m}$ can be represented by a vertically periodic $2^{k}\times(2^{n(m+1)-k}+m)$ matrix. If the $2^{k}\times m$ matrix obtained from the first $m$ columns of this matrix equals to the $2^{k}\times m$ matrix obtained from the last $m$ columns, then the cycle can be represented by a $2^{k}\times 2^{n(m+1)-k}$ doubly-periodic matrix. This is satisfied when $m>1$ or $n\neq k$ and the associated matrix forms a $(2^{k},2^{n(m+1)-k};n,m+1)$-DBA.
We are not interested in a code with one codeword as these codes are just de Bruijn arrays and it was proved by Paterson [42] that the necessary conditions of Theorem 1 are sufficient for this case. However, these codes can be used in our recursive constructions.
@@
An $(2^{k},2^{t};n,m)$-DBAC, where the columns of the arrays are taken from a PF$(n,k)$, say ${\cal F}$, with $\Delta$ codewords can be generated from $G_{{\cal F},m}$. Such a code in which all the columns of the codewords are from ${\cal F}$ can be viewed as a factor in $G_{{\cal F},m}$. In this factor of $G_{{\cal F},m}$ the associated $2^{k}\times 2^{t}$ matrices must be doubly periodic. In other words, this factor in $G_{{\cal F},m}$ with $\Delta$ cycles of the same length $2^{t}$ forms a $(2^{k},2^{t};n,m)$-DBAC. Any cycle of length $2^{t}$ in $G_{{\cal F},m}$ can be represented by a $2^{k}\times(2^{t}+m-1)$ matrix. But, to be a codeword in a DBAC this is conditioned to the requirement that when a codeword is represented by its $2^{k}\times(2^{t}+m-1)$ matrix, the $2^{k}\times(m-1)$ matrix obtained from the first $m-1$ columns and the last $m-1$ columns of each matrix are equal. It should be noted that there are $(2^{k},2^{t};n,m)$-DBACs whose columns do not form a perfect factor in $G_{n}$. In all the direct constructions which will be presented in this paper the columns of the arrays will be cycles from a perfect factor. Moreover, the recursive constructions which be presented will make usage of the cycles in a perfect factor to determine when ${\bf D}_{0}^{-1}$ will be applied on columns of the array and when ${\bf D}_{1}^{-1}$ will be applied. This view of DBACs can be considered as a two-dimensional generalization for perfect factors. Moreover, the graph $G_{{\cal F},m}$ can be viewed as a two-dimensional generalization for the de Bruijn graph and a factor in the graph as a state diagram for a two-dimensional generalization for nonsingular feedback shift registers.
@@
The generalization for two-dimensional feedback shift register using a perfect factor PF$(n,k)$ is straightforward. Assume that perfect factor ${\cal F}$ has $2^{n-k}$ cycles $X_{0},X_{1},\ldots,X_{2^{n-k}-1}$ of length $2^{k}$ and the register has $m$ cells $x_{1},x_{2},\ldots,x_{m}$. At each stage the cell $x_{\ell}$, $1\leq\ell\leq m$, stores a column vector $X_{i_{\ell}}$ of length $2^{k}$ which is a sequence of ${\cal F}$ in some shift ${\bf E}^{j_{\ell}}$ related to its zero state, where $j_{1}=0$, i.e., the sequence in $x_{1}$ is taken in its zero state. Using the feedback function $x_{m+1}=f(x)=f(x_{1},x_{2},\ldots,x_{m})$ we have that
@@
i.e., $x_{m+1}$ equals to ${\bf E}^{j_{m+1}}X_{i_{m+1}}$, i.e., for each $\ell$, $11\leq\ell\leq m-$ 1 in the next stage controlled by a master clock $x_{\ell}$ receives its information from $x_{\ell+1}$ and hence it stores ${\bf E}^{j_{\ell+1}-j_{2}}X_{i_{\ell+1}}$. The shifts are taken modulo $2^{k}$, the length of the cycles in ${\cal F}$ and they are reduced by $j_{2}$ since in $x_{1}$ the sequence should be taken in its zero state. Finally $x_{m}$ will store ${\bf E}^{j_{m+1}-j_{2}}X_{i_{m+1}}$. The state diagram is a subgraph of $G_{{\cal F},m}$ and for the feedback function $f$ it has the edge
@@
Due to this definition and the state diagram of the feedback function we call the vertices of $G_{{\cal F},m}$ also states. As in the one-dimensional case we define a nonsingular two-dimensional feedback shift register to be a feedback shift register whose state diagram contains only cycles, i.e., it is a factor of $G_{{\cal F},m}$ associated with doubly-periodic matrices. We can now develop a theory for two-dimensional feedback shift registers in the same ways as was done in the one-dimensional case. Anyway, this complete theory is beyond the scope of the current paper, although some small fractions that are relevant to the current work will be mentioned.
The merge-or-split method is one feature that can be generalized from a factor in $G_{n}$ to a factor in $G_{{\cal F},m}$, i.e., from one-dimensional feedback shift register to two-dimensional feedback shift registers, where ${\cal F}$ is a factor in $G_{n}$ with cycles of length $2^{k}$. For simplicity we will consider factors in $G_{{\cal F},m}$ in which all the cycles have the same length $2^{t}$ and each associated $2^{k}\times 2^{t}$ matrix form an $(2^{k},2^{t};n,m)$-DBAC, i.e., the $2^{k}\times 2^{t}$ matrix is doubly periodic. A set of vertices in $G_{{\cal F},m}$ represented by $2^{k}\times m$ matrices is called a companion set if the $2^{k}\times(m-1)$ matrices defined by first $m-1$ columns of the matrices in the set are all equal. Two distinct vertices from this set will be called companion vertices (states) similarly to such pair of vertices in $G_{n}$. Two cycles which contain a pair of companion vertices from such a set, one vertex on each cycle, can be combined into one cycle when the merge-or-split method is applied on this pair of companion vertices in these two cycles. The associated $2^{k}\times 2^{t+1}$ matrix obtained from this join is doubly periodic since the sum of shifts in each cycle is 0 modulo $2^{k}$ which implies that sum of shifts in the combined cycle is also 0 modulo $2^{k}$. The reason is that the shifts in the two cycles are also the shifts of the combined cycle.
III Direct Constructions for de Bruijn Arrays Codes
In this section, we present five direct constructions for de Bruijn arrays codes. These constructions can be used for the recursive constructions presented in Section IV and for the merge-or-split method applied on some DBACs in Section V, to analyze the parameters of the $(2^{k},2^{t};n,m)$-DBACs that are obtained in our work. The first construction is aimed to construct $(2^{k},2^{t};n,1)$-DBACs for all possible parameters. The key idea is to notice that in each such array the columns are sequences of length $2^{k}$ such that each $n$-tuple is contained in exactly one of these columns. This implies that if the size of the code is $\Delta$, then the union of the sequences in all the columns of all the $\Delta$ arrays form a perfect factor in $G_{n}$ and vice versa. This perfect factor has $2^{n-k}$ cycles of length $2^{k}$, i.e., it is a PF$(n,k)$. Moreover, for this perfect factor $2^{t}$ can be any divisor of $2^{n-k}$. These observations imply the following theorem.

### v2

@@ II-B The shift operator ${\bf E}$
An operator with which many of the properties of feedback shift registers can be represented is the shift operator ${\bf E}$. This operator shifts the whole sequence to the left. For a sequence ${\cal S}=s_{0}s_{1}~{}\cdots~{}s_{k-1}$ of length $k$, ${\bf E}s_{i}=s_{i+1}$, where $10\leq i<k-$ 1 and if the sequence is cyclic then also ${\bf E}s_{k-1}=s_{0}$. When the operator is applied to the cyclic sequence ${\cal S}$ we have
@@
Any periodic (cyclic) sequence satisfies a linear recursion
@@
The linear complexity $c({\cal S})$ of ${\cal S}$ is defined as the least $m$ for which (1) holds. Clearly ${c({\cal S})\leq k}$ for any sequence with period $k$, since $s_{i+k}=s_{i}$ by definition.
The sequences that are of interest in our exposition have length a power of 2. These sequences are generated by the polynomial $({\bf E}+1)^{k}$ and for each sequence ${\cal S}$ generated by this polynomial we have $({\bf E}+{\bf 1})^{k}{\cal S}=[0,0,\ldots,0]$. The following properties characterize these sequences [10].
@@
A sequence ${\cal S}$ of length $2^{n}$ has linear complexity $c$ if and only if $({\bf E}+{\bf 1})^{c-1}{\cal S}=[1^{2^{n}}]$.
@@
There are $2^{n-\lfloor\log n\rfloor-1}$ sequences which are generated by the polynomial $({\bf E}+{\bf 1})^{n+1}$ and not generated by the polynomial $({\bf E}+{\bf 1})^{n}$. All these sequences have period $2^{\lfloor\log n\rfloor+1}$ and linear complexity ${n+1}$. Each $n$-tuple is contained exactly once in one of these sequences.
We define the weight of a word $(x_{1},x_{2},\ldots,x_{n})$ as the number of nonzero entries in the word. Similarly, we define the weight of a cyclic sequence $[x_{1},x_{2},\ldots,x_{n}]$ as the number of nonzero entries in the sequence. Note, that the weight of the word $(011)$ is two, of the sequence $[011]$ is also two, but of the sequence $[011011]$ is four, although its period is three and the weight of one period is two. Other properties of sequences are as follows.
@@ II-C The ${\bf D}$-morphism
A key operator in the recursive constructions which follow is the ${\bf D}$-morphism, originally defined by Lempel [28]. The operator ${\bf D}$ is a 2-to-1 mapping from $\mathbb{F}_{2}^{n}$ on to $\mathbb{F}_{2}^{n-1}$, where $n\geq 2$. For a binary word $(x_{1},x_{2},\ldots,x_{n})$ we have
@@
By this definition we have that ${\bf D}x={\bf D}y$ if and only if $y=x$ or $y=\bar{x}$. The operator ${\bf D}$ can be applied to cyclic sequences (cycles). For a cyclic sequence $[x_{1},x_{2},\ldots,x_{n}]$ we have
@@
It is important to note that when ${\bf D}$ or ${\bf D}_{1}^{-1}$ are applied on cyclic sequences the outcome is a cyclic sequence. The definition of ${\bf D}$ implies that ${\bf D}={\bf E}+{\bf 1}$ and hence some of the results concerning the shift operator ${\bf E}$ associated with sequences of length $2^{n}$ can be given in terms of the operator ${\bf D}$, e.g.,
@@
The operator ${\bf D}$ and its inverse ${\bf D}^{-1}$ can be applied recursively on a sequence ${\cal S}$ and its derived sequences. It is defined for $k\geq 2$ as follows.
@@
The equality ${\bf D}={\bf E}+{\bf 1}$ and the properties of the operator ${\bf D}$, its inverse ${\bf D}^{-1}$, and the linear complexity imply the following.
@@
If ${\cal F}$ is a set of $\Delta$ vertex-disjoint cycles of length (period) $2^{k}$, with linear complexity smaller than $2^{k}$ (and hence of even weight), in $G_{n}$, then ${\cal D}^{-1}{\cal F}$ contains $2\Delta$ vertex-disjoint cycles of length $2^{k}$ in $G_{n+1}$.
@@
If ${\cal F}$ is a set of $\Delta$ vertex-disjoint cycles of length $2^{k}$, with linear complexity $2^{k}$ (and hence of odd weight), in $G_{n}$, then ${\cal D}^{-1}{\cal F}$ contains $\Delta$ vertex-disjoint self-dual cycles of length $2^{k+1}$ in $G_{n+1}$.
@@ II-D Perfect factors
The perfect factors of $G_{n}$, defined below, play pivotal role in this paper.
@@
The theorem was extended for larger alphabet by Paterson [44]. The following case of his result is required for Construction 4 and Theorem 6.
@@
For a given set of parameters there could be many distinct perfect factors. We will be interested in perfect factors in which the weights of all the sequences are either even or odd. Both types of factors will be important in our exposition and they will be discussed in Section VI.
@@
One type of PF$(n,n-1)$, whose sequences are ${\cal S}$ and $\bar{{\cal S}}$, is of special interest. The sequence which contains the all-zero $n$-tuple in such a factor will be called a span $n$ half de Bruijn sequence. In such a sequence either $x$ or $\bar{x}$ is contained in ${\cal S}$, for any given $n$-tuple $x$.
III Direct Constructions for de Bruijn Array Codes
In this section, we present direct constructions for de Bruijn array codes. These constructions can be used to generate input to the recursive constructions presented in Section IV and to the merge-or-split method applied on some DBACs in Section V. The section consists of three subsection. In Section III-A we present a two-dimensional generalization of the de Bruijn graph which yields also a two-dimensional generalization of feedback shift registers. Based on this generalization in Section III-B two direct constructions of DBACs are presented. Two other direct constructions for DBACs are presented in Section III-C.
III-A Two-dimensional FSR and de Bruijn graph
@@
One-dimensional sequences can be viewed as cycles in the de Bruijn graph. We aim also to view the two-dimensional arrays constructed here as cycles in the family of graphs defined now. The graph for the representation of the two-dimensional codes was defined in [8]. This graph is appropriate to codes in which all the columns of the matrices that form the codewords are sequences of some given perfect factor. For this purpose, let ${\cal F}$ be a PF$(n,k)$ and define a graph $G_{{\cal F},m}$ whose vertices are $2^{k}\times m$ matrices. These matrices are periodic (cyclic) only vertically. Each such matrix $X$ can be represented as
@@
where $X_{i_{\ell}}$, $10\leq\ell\leq m-$ 1, is a cycle in ${\cal F}$ represented as a column vector and ${\bf E}^{j_{\ell}}X_{i_{\ell}}$ is a cyclic shift of $X_{i_{\ell}}$ up by $j_{\ell}$ positions compared to its zero state, i.e., it is the column vector of ${\bf E}^{j_{\ell}}X_{i_{\ell}}$. The first cycle $X_{i_{0}}$ is taken in its zero shift, i.e., it starts with its zero state. An edge in the graph is a $2^{k}\times(m+1)$ matrix which is periodic only vertically. Such a matrix
@@
This implies that the graph $G_{{\cal F},m}$ has $2^{nm-k}$ vertices and $2^{n(m+1)-k}$ edges.
@@
A $(2^{k},2^{t};n,m)$-DBAC, where the columns of the arrays are taken from a PF$(n,k)$, say ${\cal F}$, with $\Delta$ codewords can be generated from $G_{{\cal F},m}$. A code in which all the columns of the codewords are from ${\cal F}$ can be viewed as a factor in $G_{{\cal F},m}$. In this factor of $G_{{\cal F},m}$ the associated $2^{k}\times 2^{t}$ matrices must be doubly-periodic. In other words, this factor in $G_{{\cal F},m}$ with $\Delta$ cycles of the same length $2^{t}$ forms a $(2^{k},2^{t};n,m)$-DBAC. Any cycle of length $2^{t}$ in $G_{{\cal F},m}$ can be represented by a vertically periodic (but not horizontally periodic) $2^{k}\times(2^{t}+m-1)$ matrix $A$. If the last $m-1$ columns of $A$ are the same as the first $m-1$ column (with no vertical cyclic shift), then the cycle can be represented by $2^{k}\times t$ doubly periodic matrix ${\cal C}$ and this matrix can be a codeword in a $(2^{k},2^{t};n,m)$-DBAC. It should be noted that there are $(2^{k},2^{t};n,m)$-DBACs whose columns do not form a perfect factor in $G_{n}$. In all the direct constructions presented in this paper the columns of the arrays will be cycles from a perfect factor. Moreover, the recursive constructions will make use of the cycles in a perfect factor to determine when ${\bf D}_{0}^{-1}$ is applied on columns of the array and when ${\bf D}_{1}^{-1}$ is applied. This view of DBACs can be considered as a two-dimensional generalization of perfect factors. Moreover, the graph $G_{{\cal F},m}$ can be viewed as a two-dimensional generalization of the de Bruijn graph and a factor in the graph as a state diagram for a two-dimensional generalization of nonsingular feedback shift registers.
@@
The generalization for two-dimensional feedback shift register using a perfect factor PF$(n,k)$ is straightforward. Suppose that perfect factor ${\cal F}$ has $2^{n-k}$ cycles $X_{0},X_{1},\ldots,X_{2^{n-k}-1}$ of length $2^{k}$ and the register has $m$ cells with variables $x_{1},x_{2},\ldots,x_{m}$. At each stage the cell $\ell$-th cell, $1\leq\ell\leq m$, stores a column vector $X_{i_{\ell}}$ of length $2^{k}$ which is a sequence of ${\cal F}$ in some shift ${\bf E}^{j_{\ell}}$ related to its zero state, where $j_{1}=0$, i.e., the sequence in $x_{1}$ is taken in its zero state. Using the feedback function $x_{m+1}=f(x)=f(x_{1},x_{2},\ldots,x_{m})$ we have
@@
i.e., $x_{m+1}$ equals ${\bf E}^{j_{m+1}}X_{i_{m+1}}$, i.e., for each $\ell$, $11\leq\ell\leq m-$ 1, the $\ell$-th cell receives its information from the $(\ell+1)$-th cell and hence it stores ${\bf E}^{j_{\ell+1}-j_{2}}X_{i_{\ell+1}}$. The shifts are taken modulo $2^{k}$, the length of the cycles in ${\cal F}$, and they are reduced by $j_{2}$ since in $x_{1}$ the sequence should be taken in its zero state. Finally $x_{m}$ will store ${\bf E}^{j_{m+1}-j_{2}}X_{i_{m+1}}$. The state diagram is a subgraph of $G_{{\cal F},m}$ and for the feedback function $f$ it has the edge
@@
Given this definition and the state diagram of the feedback function we also call the vertices of $G_{{\cal F},m}$ states. As in the one-dimensional case we define a nonsingular two-dimensional feedback shift register to be a feedback shift register whose state diagram contains only cycles. We can now develop a theory for two-dimensional feedback shift registers analogous to the one-dimensional case. Such theory is beyond the scope of the current paper, although some references relevant to the current work will be made.
The merge-or-split method is a feature that can be generalized from a factor in $G_{n}$ to a factor in $G_{{\cal F},m}$, i.e., from one-dimensional feedback shift register to two-dimensional feedback shift registers, where ${\cal F}$ is a factor in $G_{n}$ with cycles of length $2^{k}$. For simplicity we will consider factors in $G_{{\cal F},m}$ in which all the cycles have the same length $2^{t}$ and each associated $2^{k}\times 2^{t}$ matrix form an $(2^{k},2^{t};n,m)$-DBAC, i.e., the $2^{k}\times 2^{t}$ matrix is doubly-periodic (as $2^{k}\times(2^{t}+m-1)$ vertically-periodic array the first $m-1$ columns equal to the last $m-1$ columns). A set of vertices in $G_{{\cal F},m}$ represented by $2^{k}\times m$ matrices is called a companion set if the $2^{k}\times(m-1)$ matrices defined by first $m-1$ columns of the matrices in the set are all equal. Two distinct vertices from this set will be called companion vertices (states) analogously to such pair of vertices in $G_{n}$. Two cycles which contain a pair of companion vertices from such a set, one vertex on each cycle, can be combined into one cycle when the merge-or-split method is applied to this pair of companion vertices in these two cycles. The associated $2^{k}\times 2^{t+1}$ matrix obtained from this join is doubly-periodic since the sum of shifts in each cycle is 0 modulo $2^{k}$ which implies that sum of shifts in the combined cycle is also 0 modulo $2^{k}$. The reason is that the shifts in the two cycles are also the shifts of the combined cycle.
III-B Codes from two-dimensional FSRs
The first two direct constructions are based on the existence of some perfect factors, and both can be viewed as a two-dimensional generalization of nonsingular FSR.
Construction 1
Let ${\cal F}$ be a PF$(n,k)$, $m=2^{t}-1$, $k\leq t$, $2\leq k<n<2^{k}$, and define the following two-dimensional FSR
$f(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{m}}X_{i_{m}})={\bf E}^{j_{m+1}}X_{i_{m+1}},$ (3)
where
$0\leq j_{r}\leq 2^{k}-1,~{}2\leq r\leq m+1\},$
where each $X_{i_{r}}$ is a column vector and its cyclic shift ${\bf E}^{j_{r}}$ is relative to its zero state. Let $\mathbb{C}$ be the code defined by the cycles of $f$.